Ask
your own question, for FREE!
Probability
9 Online
The probability of contamination in batch 1 of a drug (event A) is 0.16, and the probability of contamination in batch 2 of the drug (event B) is 0.09. The probability of contamination in batch 2, given that there was a contamination in batch 1, is 0.12. Given this information, which statement is true?
Still Need Help?
Join the QuestionCove community and study together with friends!
|dw:1398603966889:dw| The flow of question can be seen by drawing a simple diagram. For event B occurring after event A occurred, the probability (follow the arrows) is 0.16x0.09=0.0144
Events A and B are not independent because P(B|A) ≠ P(B).
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
6 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal